OpenStudy (anonymous):

A ball is thrown straight down from the top of a 304 ft building with an initial velocity of -12 feet per second. The position function is s(t)= -16t^2-12t+304. What is its velocity after 2 seconds? What is its velocity after falling 180 feet?

OpenStudy (anonymous):

\[v _{0}=-12 fps\]g=-32fps \[v=v _{0}-g*t\] v=-12-(-32*2) v=-76fps Find the time at 180 feet \[180=-16t ^{2}-12t+304\]\[0=-16t ^{2}-12t+124\]Solve for t using the quadratic equation. Once you find t, use the velocity equation above using the new time.